OpenStudy (anonymous):

Solve the following problem. It may be helpful to use draw a chart on scrap paper to organize the information and write the equation. Be sure to show all steps (V.E.S.T.) and work in order to receive full credit. A grocer wants to make a 10-pound mixture of cashews and peanuts that he can sell for $3.29 per pound. If cashews cost $5.60 per pound and peanuts cost $2.30 per pound, how many pounds of each must he mix? help /.\

OpenStudy (♪chibiterasu):

Using 2 equations: Quantity: c + p = 10 lbs Value::: 5.6c + 2.3p = 3.29*10 dollars ---------- Multiply thru the Quantity Eq. by 560 Multiply thru the Value Eq. by 100 ------------------- 560c + 560p = 5600 560c + 230p = 3290 ----- Subtract and solve for "p": 330p = 2310 p = 7 lbs (amt. of peanuts needed) ------ Solve for "c" using c+p = 10 c + 7 = 10 c = 3 lbs (amt. of cashews needed)

OpenStudy (anonymous):

thank you sooo much :)

OpenStudy (♪chibiterasu):

Glad to help c:

OpenStudy (anonymous):

^.\